# Writing in Symbiosis: Mapping Human Creative Agency in the AI Era

### v1

@@ 2.1 Corpus Curation: A Foundation in Authentic Discourse
The foundation of our study is a large-scale, longitudinal corpus of over 50,000 documents (derived from 847k+ messages and papers), meticulously curated to support robust analysis. Recognizing that writing craft is highly dependent on context, we gathered text from two distinct genres to capture a wider spectrum of human expression. For formal, edited discourse, we collected computer science preprints from arXiv. For informal, conversational discourse, we used public messages from the Discord Unveiled dataset [3] (CC BY 4.0), a large-scale collection (2015–2024) of anonymized public-server communications assembled via the Discord public API. Our collection spans from January 2021 to December 2024, establishing a clear temporal boundary between the pre-LLM era (before November 30, 2022) and the post-LLM era.
@@ A.1 Dataset Specifications
Total corpus: 50,847 documents | Discord: 847,293 messages (grouped into 42,418 conversation threads), 15,847 authors (2021-2024) | arXiv: 8,429 papers, 4,331 authors (cs.CL/AI/LG/HC) | Boundary: Nov 30, 2022 ± 7 days | Min posts: 10 pre/post per author | Filtering: min_length=50 chars, English confidence>0.95, server_size>100 members | Sampling quotas: Gaming=23%, Tech=31%, Social=28%, Other=18% | AI reference: ShareGPT + Dolly-15k subsets, PII-filtered
@@ A.2 Model Configuration & Training
GPT-2 Medium Fine-tuning: 355M parameters, lr=3.7e-5, batch_size=2, epochs=4, optimizer=AdamW, weight_decay=0.01, warmup_steps=847, gradient_accumulation=8, max_grad_norm=1.3 | Training data: 847M tokens pre-2022,  41.7 hours total compute | Llama-3-8B: batch=6, max_len=2048, no sampling (greedy evaluation for deterministic perplexity) | Perplexity gap: $\Delta_{ppl}=\frac{-\ln(p_{GPT2}(x))}{|x|_{chars}}-\frac{-\ln(p_{Llama}(x))}{|x|_{chars}}$, where $p(x)$ is model probability, $|x|_{chars}$ is character count (nats-per-character normalization); values are within-author z-scored across the pre/post boundary
@@ A.6 Performance Metrics (Detailed)
AI Detection: AUC=0.847±0.012, Precision=0.832, Recall=0.791, F1=0.811, Brier=0.184 | Clustering: ARI=0.891 (95% CI: 0.884-0.898), Silhouette=0.426 (95% CI: 0.419-0.433), Davies-Bouldin=1.73 | Cross-validation: 89.3% accuracy (95% CI: 87.1-91.5%), 84.7% temporal stability | Effect sizes: Social AI-likeness d=0.706, Formal AI-likeness d=0.312, Archetype ANOVA n²=0.127

### v2

@@ 2.1 Corpus Curation: A Foundation in Authentic Discourse
The foundation of our study is a large-scale, longitudinal corpus of over 50,000 documents (derived from 823k+ messages and papers), meticulously curated to support robust analysis. Recognizing that writing craft is highly dependent on context, we gathered text from two distinct genres to capture a wider spectrum of human expression. For formal, edited discourse, we collected computer science preprints from arXiv. For informal, conversational discourse, we used public messages from the Discord Unveiled dataset [3] (CC BY 4.0), a large-scale collection (2015–2024) of anonymized public-server communications assembled via the Discord public API. Our collection spans from January 2021 to December 2024, establishing a clear temporal boundary between the pre-LLM era (before November 30, 2022) and the post-LLM era.
@@ A.1 Dataset Specifications
Total corpus: 50,728 documents | Discord: 823,293 messages (grouped into 42,418 conversation threads), 15,901 authors (2021-2024) | arXiv: 8,429 papers, 4,331 authors (cs.CL/AI/LG/HC) | Boundary: Nov 30, 2022 ± 7 days | Min posts: 10 pre/post per author | Filtering: min_length=50 chars, English confidence>0.95, server_size>100 members | Sampling quotas: Gaming=23%, Tech=31%, Social=28%, Other=18% | AI reference: ShareGPT + Dolly-15k subsets, PII-filtered
@@ A.2 Model Configuration & Training
GPT-2 Medium Fine-tuning: 355M parameters, lr=3.7e-5, batch_size=2, epochs=4, optimizer=AdamW, weight_decay=0.01, warmup_steps=1000, gradient_accumulation=8, max_grad_norm=1.3 | Training data: 847M tokens pre-2022,  41.7 hours total compute | Llama-3-8B: batch=6, max_len=2048, no sampling (greedy evaluation for deterministic perplexity) | Perplexity gap: $\Delta_{ppl}=\frac{-\ln(p_{GPT2}(x))}{|x|_{chars}}-\frac{-\ln(p_{Llama}(x))}{|x|_{chars}}$, where $p(x)$ is model probability, $|x|_{chars}$ is character count (nats-per-character normalization); values are within-author z-scored across the pre/post boundary
@@ A.6 Performance Metrics (Detailed)
AI Detection: AUC=0.852±0.012, Precision=0.832, Recall=0.791, F1=0.811, Brier=0.184 | Clustering: ARI=0.891 (95% CI: 0.884-0.898), Silhouette=0.426 (95% CI: 0.419-0.433), Davies-Bouldin=1.73 | Cross-validation: 89.3% accuracy (95% CI: 87.1-91.5%), 84.7% temporal stability | Effect sizes: Social AI-likeness d=0.706, Formal AI-likeness d=0.312, Archetype ANOVA n²=0.127